General Hospital , a not-for-profit acute care facility, has estimated the following costs for its inpatient services: Fixed Costs $ 15,000,000 Variable Cost Per Inpatient day $600 The hospital expects to have 8,000 inpatient days next year. a. What is the hospital’s underlying cost structure? b. What are the hospitals expected total cost? c. What is the hospital’s estimated total cost at 12,500 inpatient days? At 17,500patient days?
Answer #1

a). hospital's underlying cost structure = Fixed cost + variable cost*no. of inpatient days

= 15000000 + (600*V)

b). hospital’s expected total cost

TC = 15000000+ (600*8000) = 15000000 + 4800000 = $19800000

c). hospital’s estimated total cost at 12500 Inpatient days

TC = 15000000 + (600*12500) = $22500000

at 17500 inpatient days

TC = 15000000+ (600*17500) = $25500000